# Performance and radiation damage mitigation strategy for silicon photomultipliers on LEO space missions

### v2

@@ 1 Introduction: silicon photosensors in space applications
The NUSES [5] project will employ SiPMs for both of its payloads, Terzina [6] and Ziré [7]. Ziré, dedicated to monitoring of low-energy ($<250$ MeV) cosmic ray fluxes in the Van Allen belts, gamma-ray bursts in the energy region 1-10 MeV, space weather and lithosphere-ionosphere- magnetosphere couplings, will employ SiPMs with LYSO/GAGG crystals in the calorimeter (CALOg), in the plastic scintillator tower (PST) reading scintillating tiles, and in the FTK fiber tracker.
@@
Several other space missions are leveraging SiPM technology to advance our understanding of high-energy astrophysical phenomena. SiPMs are foreseen to be used for the HERD (High Energy Cosmic Radiation Detection) plastic scintillator detector [9]. HERD is a planned space-based experiment to study dark matter from space through the measurement of energy spectra and anisotropy of high energy electrons, as gamma-ray monitoring between 0.5 GeV and 100 TeV and to measure cosmic ray fluxes to the knee. A future mission following the paths of Ziré employing scintillating crystals read by SiPMs is Crystal Eye [10].
SiPMs play a crucial role in Cherenkov light detection in Terzina. The proposed POEMMA (Probe Of Extreme Multi-Messenger Astrophysics) mission [11, 12] firstly proposed to complement measurements of fluorescence with MAPMTs with Cherenkov detection with the more sensitive SiPMs. Cherenkov light is emitted by ultra-high-energy cosmic rays (UHECRs) and can also be detected from earth skimming astrophysical neutrinos when they produce extensive air showers (EAS) in the Earth’s atmosphere. This mission is preceded by several balloon flights under the JEM-EUSO (Joint Experiment Missions for Extreme Universe Space Observations) program [13, 14]. In particular, EUSO-SPB2 (Super Pressure Balloon-borned Telescope) [15, 16], launched in May 2023 from NASA’s Wanaka base in New Zealand [17], employed SiPMs for cosmic-ray detection from the stratosphere.
This study for Terzina, where the photosensor plane is directly exposed to cosmic radiation, is imperative. A previous study was performed by POLAR-2 [18], where SiPMs are coupled with scintillators as for other experiments. They amplify weak signals and provide protection against radiation damage, along with the shielding provided by the satellite structure. The POLAR-2 mission plans to use SiPMs for advanced gamma-ray burst (GRB) polarization studies and radiation hardness studies. Specifically regarding radiation hardness, they have proposed mitigation strategies that align with the work described here [19, 18].
@@ 2 The Terzina telescope onboard NUSES
The NUSES ballistic mission aims to explore new scientific and technological pathways for future space-based astroparticle physics detectors. The satellite platform, developed by the industrial partner TAS-I (THALES Alenia Space - Italia), will operate in a quasi-polar Low Earth Orbit (LEO) at an altitude of 535 in the worst case and 550 km at the Beginning of Life (BoL). The satellite will host two instruments: Terzina and Ziré. It will follow a sun-synchronous orbit and operate in a dusk-dawn mode along the day/night boundary so that the Terzina telescope can point to the dark side of the Earth’s limb to detect the faint Cherenkov light from ultra-relativistic particle showers in the dark atmosphere. Terzina telescope aims at the first detection of this Cherenkov light, which can be detected within the field of view (FoV) of its camera located in the FPA. The FPA is designed to detect the Cherenkov light emitted by UHECR-induced EAS with energies above $\sim 100$ PeV and with their axis pointing upwards close to the optical axis of the telescope pointing to the limb.
@@
The Terzina detector consists of a near-UV optical telescope with Schmidt-Cassegrain optics (shown in Fig. 3 with effective collective area of 0.09 m2 and equivalent focal length of 925 mm and a FPA in the focal plane. The shadowing due to the support of the system (baffles and vanes) is 8%. The Optical Telescope Assembly (OTA), which will be realised by the company Officina Stellare, is composed of a hyperbolic primary mirror of 430 mm diameter and radius of curvature (RoC) of 1200 mm, fitting a flat FPA of 131$\times$60 mm2. The secondary aspherical mirror has diameter of 194 mm and RoC = 605 mm and it is located at a distance of 430 mm from the primary mirror. The secondary mirror is covered by a correcting lens of thickness 30 mm and RoC = 320 mm to compensate for all aberrations and principally for those induced by the flat FPA (if curved, it would have reduced aberrations but the mechanical design of the FPA would have been more complex).
@@ 3 The NUV-HD-MT SiPMs for Terzina
    cross-talk of less than 10% at the operation voltage;
@@
In partnership with FBK, we have selected their technology minimizing the cross-talk and the DCR, the NUV-HD-MT [25]. We then investigated the micro-cell size that maximize the PDF while keeping the signal duration short. In fact, the signal duration is proportional to the micro-cell size and its capacitance. A large signal increases the noise captured within the acquisition window required to integrate the full pulse. The NUV-HD-MT technology has been developed by FBK adding metal-filled Deep Trench Isolation (DTI) to strongly suppress optical cross-talk. Compared to NUV-HD, FBK finds a reduction in cross-talk by a factor of 10 [26]. We have measured the NUV-HD-MT SiPM of $3\times 3$ mm2 and $1\times 1$ mm2 with square micro-cell sizes of 25, 30, 35, 40, 50 $\mu\text{m}$. A further significant consideration is whether to utilise bare or coated sensors. While bare sensors require meticulous handling, the electrons may radiate Cherenkov light within the covering resin layer (see section 8).
@@ 3.1 Static characterisation
The static characterisation and the forward IV measurements have been conducted to estimate the quenching resistance, while the reverse IV measurements have been used to determine the breakdown voltage.
@@
  • Pulsed LED with 450 nm wavelength, powered by a pulse generator operating at the frequency of 10 kHz. This creates alternating intervals: one where the device operates in complete absence of light and another where it is illuminated by the LED light (synchronized with the trigger waveform acquisition), during which photon signals are expected. Hereafter, these time intervals are referred to as dark interval and trigger window, respectively.
  • Integration sphere, used to eliminate any spatial variations in the incoming light by multiple scattering reflections on its internal surface. This ensures a uniform diffuse light output and preserves the power at each output port.
  • Neutral density filter, is inserted between the integration sphere output port and SiPM. For an automated switch between each filter, it is mounted on a motorized wheel. This enables us to illuminate the SiPMs with different light intensities as needed.
  • Diffuser, is inserted right after the neutral density filter to further reduce any residual non-uniformity from the output port of the integration sphere inside the collimator leading to the sensors.
  • Control photodiode, it is mounted at another output port of the integration sphere to monitor the amount of the inserted light within the setup at all times. This is connected to a Keithley 6487 picoammeter [^1] The Keithley 6487 picoammeter is able to perform measurements of the current with precision of around 6.4 pA for currents lower than 2 nA./voltage source [28].
  • Temperature and humidity sensor, inserted near the SiPM and monitors the temperature constantly during the measurement.
  • Analyser PC, which monitors and logs the output from all measurement devices.
For the static characterization configuration, the IV curve was measured by directly biasing and reading the SiPM using a Keithley 6487 or Keithley 2400 [29] picoammeter/voltage source. In the dynamic characterization configuration, the bias voltage was supplied by a Keithley 6487 picoammeter/voltage source reading before the signal with a 2 GHz bandwidth oscilloscope. A custom-made amplifier with a bandwidth of approximately $\sim 600$ MHz and a gain of $\sim 40$ dB was used to amplify, pulse shape and filter the SiPM signal, and effectively suppressing the slow component of the SiPM response. The Keithley 2000 multimeter [30] shown in this configuration is used for reading a temperature sensor mounted near the control photodiode. The average temperature during the measurements was $T=22$ C (hereinafter referred to as room temperature).
@@ 3.1.1 Forward IV characterisation
Figure 5: The forward (left) and reverse (right) IV characteristic data points with scale errors of the instruments for the coated FBK SiPM with 30 $\mu\mathrm{m}$ micro-cell size. The errors are due to the instrument scale precision.
@@ 3.1.2 Reverse IV characterisation and IV model for post-breakdown
In the reverse IV curves, two distinct regimes can be observed: the pre-breakdown “linear” regime and the post-breakdown Geiger regime, as illustrated in figure 5. This work employs the so-called “IV model” [32, 27], which describes the behaviour of SiPMs over their complete operating range, including the Geiger regime. The IV curves for all temperatures and devices exhibit similar behaviour as it is shown in figure 6, indicating multiple regimes of bias voltage. Taking all contributions into account, the post-breakdown current, $I_{\mathrm{post-BD}}$, can be expressed as [^2] The formula does not include the pre-breakdown term.:
@@
$1\times 1$ mm2 25 $\mu$m coated $0.47\pm 0.30$ $60.0$ $32.6$ $4.21038$ $-3.33939\cdot 10^{-1}$
$1\times 1$ mm2 30 $\mu$m coated $1.00\pm 0.37$ $60.0$ $32.6$ $4.59777$ $-3.55651\cdot 10^{-1}$
$1\times 1$ mm2 35 $\mu$m coated $1.26\pm 0.39$ $60.0$ $32.6$ $4.59028$ $-5.94936\cdot 10^{-1}$
$1\times 1$ mm2 50 $\mu$m coated $1.93\pm 0.55$ $60.0$ $32.6$ $4.88520$ $-4.26954\cdot 10^{-1}$
$1\times 1$ mm2 25 $\mu$m bare $0.54\pm 0.21$ $60.0$ $32.6$ $2.85255$ $-4.06597\cdot 10^{-1}$
$1\times 1$ mm2 30 $\mu$m bare $0.59\pm 0.40$ $60.0$ $32.6$ $3.08427$ $-4.32636\cdot 10^{-1}$
$1\times 1$ mm2 35 $\mu$m bare $1.22\pm 0.61$ $60.0$ $32.6$ $3.77057$ $-4.33696\cdot 10^{-1}$
$1\times 1$ mm2 50 $\mu$m bare $2.03\pm 0.58$ $60.0$ $32.6$ $4.31814$ $-3.80479\cdot 10^{-1}$
$3\times 3$ mm2 40 $\mu$m bare $16.9\pm 4.3$ $60.0$ $32.6$ $3.68959$ $-3.41690\cdot 10^{-1}$
$3\times 3$ mm2 40 $\mu$m coated $17.8\pm 4.5$ $60.0$ $32.6$ $3.85838$ $-3.24583\cdot 10^{-1}$
@@ 3.2 Dynamic characterisation
    This noise is independent of the light conditions. At room temperature, the DCR is predominantly caused by the thermal generation of carriers. These carriers, under the operational voltage ($V_{\text{bias}}$), and influenced by the trap-assisted tunnelling effect, transition from the valence band to the conduction band, resulting in spurious signals. The increase in the DCR as a function of $V_{\text{bias}}$, due to the heightened electrical field, can be modelled by the empirical formula:
@@
    This noise arises from secondary avalanches occurring within the same or neighbouring micro-cells and can be categorized into:
@@
To perform the dynamic characterisation of SiPM and measure the uncorrelated noise, we utilized the setup shown in the right panel of Figure 4. A full over-voltage scan was conducted by ranging the voltage betweem 1 V and 18 V, and by acquiring 10’000 waveforms on the oscilloscope for each given over-voltage. To compute the DCR and OCT, the analysis specifically focused on the signals received within dark interval of the waveform (in the absence of light).
@@
Figure 9: a) Our measurements of DCR per mm2 as a function of the over-voltage at room temperature for sensors without coating with different micro-cell sizes. The measurements of FBK are indicated in the label. b) Our and FBK measurements of OCT as a function of over-voltage for sensors of 25 and 30 $\mu$m micro-cell sizes without coating at room temperature.
@@ 3.3 Optical characterisation
Weighing the DCR and PDE, as described above, against the pulse shape as outlined in [8], it has been decided to use SiPM tiles with a 30 $\mu$m bare $\mu$-cell size. At this configuration, we have an optimal performance in DCR and PDE, whilst still having a short pulse decay time. The decay time and the PDE are anti-correlated: smaller $\mu$-cell sizes result in shorter signal durations but also in lower PDE due to the reduced fill factor. Hence, the final choice of the SiPM $\mu$-cell of 30 $\mu$m represents a trade-off between these parameters. Additionally it complies with the constraints on the maximum power consumption of the NUSES satellite. As a matter of fact, the larger the DCR, the larger the power consumption of the SiPM power supplies. If the data acquisition thresholds are not adjusted consequently, the rate of noise induced triggers would also require more processing activity further contributing to the increase in power consumption. We performed a set of measurements with three different configurations: 1) with a pulsed LED (450 nm) and an integration sphere; 2) with a short-duration ($\sim$25 ps) light-pulse laser with a 370 nm wavelength laser with and 3) without the integration sphere. Both the LED and the laser flash the full sensor to obtain a high amplitude signal to be detectable by a 2 GHz bandwidth oscilloscope. Figure 11 summarizes the obtained results. As expected, the shortest decay time was measured with the laser only, as the integration sphere induces an additional time spread. The SiPM signal tail has been fitted with a negative exponential function and the decay time is $\tau$ in the figure. This time is a function of the $\mu$-cell capacitance, hence its size. We did not observe any significant change in the signal shape with the variation in the bias voltage.

### v3

@@ 1 Introduction: silicon photosensors in space applications
The NUSES [5] project will employ SiPMs for both of its payloads, Terzina [6] and Ziré [7]. Ziré, dedicated to monitoring of low-energy ($<250$ MeV) cosmic ray fluxes in the Van Allen belts, gamma-ray bursts in the energy region 1-10 MeV, space weather and lithosphere-ionosphere- magnetosphere couplings, will employ SiPMs with LYSO or GAGG crystals in the calorimeter (CALOg), in the plastic scintillator tower (PST) reading scintillating tiles, and in the FTK fiber tracker.
@@
Several other space missions are leveraging SiPM technology to advance our understanding of high-energy astrophysical phenomena. SiPMs (produced by Hamamatsu Photonics) are foreseen to be used for the HERD (High Energy Cosmic Radiation Detection) plastic scintillator detector [9]. HERD is a planned space-based experiment to study dark matter from space through the measurement of energy spectra and anisotropy of high energy electrons, as gamma-ray monitoring between 0.5 GeV and 100 TeV and to measure cosmic ray fluxes to the knee. A future mission following the paths of Ziré employing scintillating crystals read by SiPMs is Crystal Eye [10].
SiPMs play a crucial role in Cherenkov light detection in Terzina. The proposed POEMMA (Probe Of Extreme Multi-Messenger Astrophysics) mission [11, 12] firstly proposed to complement measurements of fluorescence with MAPMTs with Cherenkov detection with the more sensitive SiPMs. Cherenkov light is emitted by ultra-high-energy cosmic rays (UHECRs) and can also be detected from earth-skimming astrophysical neutrinos when they produce extensive air showers (EAS) in the Earth’s atmosphere. This mission is preceded by several balloon flights under the JEM-EUSO (Joint Experiment Missions for Extreme Universe Space Observations) program [13, 14]. In particular, EUSO-SPB2 (Extreme Universe Space Observatory on a Super Pressure Balloon 2) [15, 16], launched in May 2023 from NASA’s Wanaka base in New Zealand [17], employed SiPMs, produced by Hamamatsu Photonics, for cosmic-ray detection from the stratosphere.
This study for Terzina, where the photosensor plane is directly exposed to cosmic radiation, is imperative. A previous study was performed on S13361-6075NE-04 and S14161-6050HS-04 from Hamamatsu by POLAR-2 [18], where SiPMs are coupled with scintillators as for other experiments. They amplify weak signals and provide protection against radiation damage, along with the shielding provided by the satellite structure. The POLAR-2 mission plans to use SiPMs for advanced gamma-ray burst (GRB) polarization studies and radiation hardness studies. Specifically regarding radiation hardness, they have proposed mitigation strategies that align with the work described here [19, 18].
@@ 2 The Terzina telescope onboard NUSES
The NUSES ballistic mission aims to explore new scientific and technological pathways for future space-based astroparticle physics detectors. The satellite platform, developed by the industrial partner TAS-I (THALES Alenia Space - Italia), will operate in a quasi-polar Low Earth Orbit (LEO) at an altitude of 535 in the worst case and 550 km at the Beginning of Life (BoL). The satellite will host two instruments: Terzina and Ziré. It will follow a sun-synchronous orbit and operate in a dusk-dawn mode along the day/night boundary so that the Terzina telescope can point to the dark side of the Earth’s limb to detect the faint Cherenkov light from ultra-relativistic particle showers in the dark atmosphere. The Terzina telescope aims at the first detection of this Cherenkov light, which can be detected within the field of view (FoV) of its camera located in the FPA. The FPA is designed to detect the Cherenkov light emitted by UHECR-induced EAS with energies above $\sim 100$ PeV and with their axis pointing upwards close to the optical axis of the telescope pointing to the limb.
@@
The Terzina detector consists of a near-UV optical telescope with Schmidt-Cassegrain optics (shown in figure 3 with effective collective area of 0.09 m2 and equivalent focal length of 925 mm and a FPA in the focal plane. The shadowing due to the support of the system (baffles and vanes) is 8%. The Optical Telescope Assembly (OTA), which will be realised by the company Officina Stellare, is composed of a hyperbolic primary mirror of 430 mm diameter and radius of curvature (RoC) of 1200 mm, fitting a flat FPA of 131$\times$60 mm2. The secondary aspherical mirror has diameter of 194 mm and RoC = 605 mm and it is located at a distance of 430 mm from the primary mirror. The secondary mirror is covered by a correcting lens of thickness 30 mm and RoC = 320 mm to compensate for all aberrations and principally for those induced by the flat FPA (if curved, it would have reduced aberrations but the mechanical design of the FPA would have been more complex).
@@ 3 The NUV-HD-MT SiPMs for Terzina
    optical cross-talk of less than 10% at the operation voltage;
@@
In partnership with FBK, we have selected their technology minimizing the optical cross-talk and the DCR, the NUV-HD-MT [25]. We then investigated the micro-cell size that maximize the PDE while keeping the signal duration short. In fact, the signal duration is proportional to the micro-cell size and its capacitance. A large signal increases the noise captured within the acquisition window required to integrate the full pulse. The NUV-HD-MT technology has been developed by FBK adding metal-filled Deep Trench Isolation (DTI) to strongly suppress optical cross-talk. FBK finds a reduction in optical cross-talk by a factor of 10 [26] with respect to the NUV-HD in absence of the metall-filled DTI. We have measured the NUV-HD-MT SiPM of $3\times 3$ mm2 and $1\times 1$ mm2 with square micro-cell sizes of 25, 30, 35, 40, 50 $\mu\text{m}$. A further significant consideration is whether to utilise bare or coated sensors. While bare sensors require meticulous handling, the electrons may radiate Cherenkov light within the covering resin layer (see section 8).
@@ 3.1 Static characterisation
The static characterisation has been done: the forward IV measurements have been conducted to estimate the quenching resistance, while the reverse IV measurements have been used to determine the breakdown voltage.
@@
  • a pulsed LED with 450 nm wavelength, powered by a pulse generator operating at the frequency of 10 kHz. This creates alternating intervals: one where the device operates in complete absence of light and another where it is illuminated by the LED light (synchronized with the trigger waveform acquisition), during which photon signals are expected. Hereafter, these time intervals are referred to as dark interval and trigger window, respectively;
  • an integration sphere which eliminates any spatial variations in the incoming light by multiple scattered reflections on its internal surface. This ensures a uniform diffuse light output and preserves the power at each output port;
  • a neutral density filter inserted between the integration sphere output port and the SiPM. It is mounted on a motorised wheel for automatic switching between each filter. This allows us to illuminate the SiPMs with different light intensities as required;
  • a diffuser, inserted immediately after the neutral density filter, to further reduce any residual non-uniformity from the output port of the integration sphere inside the collimator to the sensors;
  • a control photodiode, mounted on another output port of the integration sphere to monitor the amount of the light within the setup at all times; this is connected to a Keithley 6487 picoammeter [^1] The Keithley 6487 picoammeter is able to perform measurements of the current with precision of around 6.4 pA for currents lower than 2 nA./voltage source [28];
  • a temperature and humidity sensor, placed near the SiPM, which constantly monitors the temperature during the measurement;
  • An analyser PC that monitors and records the output of all instruments.
For the static characterization configuration, the IV curve was measured by directly biasing and reading the SiPM using a Keithley 6487 or Keithley 2400 [29] picoammeter/voltage source. In the dynamic characterization configuration, the bias voltage was supplied by a Keithley 6487 picoammeter/voltage source reading before the signal with a 2 GHz bandwidth oscilloscope. A custom-made amplifier with a bandwidth of approximately $\sim 600$ MHz and a gain of $\sim 40$ dB was used to amplify, pulse shape and filter the SiPM signal, and effectively suppressing the slow component of the SiPM response. The Keithley 2000 multimeter [30], shown in this configuration, has been used for reading a temperature sensor mounted near the control photodiode. The average temperature during the measurements was $T=22$ C (hereinafter referred to as room temperature).
@@ 3.1.1 Forward IV characterisation
Figure 5: The forward (left) and reverse (right) IV characteristic data points with scale errors of the instruments for the FBK SiPM with 30 $\mu\mathrm{m}$ micro-cell size. The errors are due to the instrument scale precision.
@@ 3.1.2 Reverse IV characterisation and IV model for post-breakdown
In the reverse IV curves, two distinct regimes can be observed: the pre-breakdown “linear” regime and the post-breakdown Geiger regime, as illustrated in figure 5. This work employs the so-called “IV model” [32, 27], which describes the behaviour of SiPMs over their complete operating range, including the Geiger regime. The IV curves for all temperatures and devices exhibit similar behaviour, as it is shown in figure 6, indicating multiple regimes of bias voltage. Taking all contributions into account, the post-breakdown current, $I_{\mathrm{post-BD}}$, can be expressed as [^2] The formula does not include the pre-breakdown term.:
@@
$1\times 1$ mm2 25 $\mu$m coated $0.47\pm 0.30$ $60.0$ $32.6$ $4.21\pm 0.86$ $(-3.34\pm 0.32)\cdot 10^{-1}$
$1\times 1$ mm2 30 $\mu$m coated $1.00\pm 0.37$ $60.0$ $32.6$ $4.60\pm 0.89$ $(-3.56\pm 0.66)\cdot 10^{-1}$
$1\times 1$ mm2 35 $\mu$m coated $1.26\pm 0.39$ $60.0$ $32.6$ $4.59\pm 0.86$ $(-5.95\pm 0.99)\cdot 10^{-1}$
$1\times 1$ mm2 50 $\mu$m coated $1.93\pm 0.55$ $60.0$ $32.6$ $4.88\pm 0.85$ $(-4.27\pm 0.54)\cdot 10^{-1}$
$1\times 1$ mm2 25 $\mu$m bare $0.54\pm 0.21$ $60.0$ $32.6$ $2.85\pm 0.95$ $(-4.07\pm 0.98)\cdot 10^{-1}$
$1\times 1$ mm2 30 $\mu$m bare $0.59\pm 0.40$ $60.0$ $32.6$ $3.08\pm 0.98$ $(-4.33\pm 0.99)\cdot 10^{-1}$
$1\times 1$ mm2 35 $\mu$m bare $1.22\pm 0.61$ $60.0$ $32.6$ $3.77\pm 0.98$ $(-4.34\pm 0.99)\cdot 10^{-1}$
$1\times 1$ mm2 50 $\mu$m bare $2.03\pm 0.58$ $60.0$ $32.6$ $4.32\pm 0.97$ $(-3.81\pm 0.47)\cdot 10^{-1}$
$3\times 3$ mm2 40 $\mu$m bare $16.9\pm 4.3$ $60.0$ $32.6$ $3.69\pm 0.91$ $(-3.42\pm 0.33)\cdot 10^{-1}$
$3\times 3$ mm2 40 $\mu$m coated $17.8\pm 4.5$ $60.0$ $32.6$ $3.86\pm 0.94$ $(-3.25\pm 0.32)\cdot 10^{-1}$
@@ 3.2 Dynamic characterisation
    This noise is independent of the light conditions. At room temperature, DCR is mainly caused by the thermal generation of carriers. Under the operation voltage ($V_{\text{bias}}$) and influenced by the trap-assisted tunneling effect, these carriers transition from the valence band to the conduction band, resulting in spurious signals. The increase in the DCR as a function of $V_{\text{bias}}$, due to the increased electric field, can be modelled by the empirical formula:
@@
    This noise arises from secondary avalanches occurring within the same or neighbouring micro-cells and can be categorised into:
@@
To perform the dynamic characterisation of SiPM and measure the uncorrelated noise, we utilized the setup shown in the right panel of Figure 4. A full over-voltage scan was conducted by ranging the voltage between 1 V and 18 V, and by acquiring 10’000 waveforms on the oscilloscope for each given over-voltage. To compute the DCR and OCT, the analysis specifically focused on the signals received within dark interval of the waveform (in the absence of light).
@@
Figure 9: a) Characterisation measurements of DCR per mm2 as a function of the over-voltage at room temperature for sensors without coating with different micro-cell sizes compared to FBK’s measurements. b) Our and FBK measurements of OCT as a function of over-voltage for sensors of 25 and 30 $\mu$m micro-cell sizes without coating at room temperature.
@@ 3.3 Optical characterisation
Weighing the DCR and PDE, as described above, against the pulse shape as outlined in [8], it has been decided to use SiPM tiles with a 30 $\mu$m bare $\mu$-cell size. At this configuration, we have an optimal performance in DCR and PDE, whilst still having a short pulse decay time. The decay time and the PDE are anti-correlated: smaller $\mu$-cell sizes result in shorter signal durations but also in lower PDE due to the reduced fill factor. Hence, the final choice of the SiPM $\mu$-cell of 30 $\mu$m represents a trade-off between these parameters. Additionally it complies with the constraints on the maximum power consumption of the NUSES satellite. As a matter of fact, the larger the DCR, the larger the power consumption of the SiPM power supplies. If the data acquisition thresholds are not adjusted accordingly, the rate of noise induced triggers would also require more processing activity, further increasing power consumption. We performed a set of measurements with three different configurations: 1) with a pulsed LED (450 nm) and an integration sphere; 2) with a short-duration ($\sim$25 ps) light-pulse laser with a 370 nm wavelength laser with and 3) without the integration sphere. Both the LED and the laser flash the full sensor to obtain a high amplitude signal to be detectable by a 2 GHz bandwidth oscilloscope. Figure 11 summarizes the obtained results. As expected, the shortest decay time was measured with the laser only, as the integration sphere induces an additional time spread. The SiPM signal tail has been fitted with a negative exponential function and the decay time is $\tau$ in the figure. This time is a function of the $\mu$-cell capacitance, hence its size. We did not observe any significant change in the signal shape with the variation in the bias voltage.